# The JADES Origins Field: A New JWST Deep Field in the JADES Second NIRCam Data Release

### v1

@@ 6 Discussion
We have described the executed and approved plans for JWST observations of the JADES Origins Field, spanning the first three mission cycles. In total, the time to be invested in this single NIRCam footprint of 9 arcmin is already very large, roughly 360 open-shutter hours of exposure time across in 15 NIRCam imaging filters, 20 hours of NIRCam slitless spectroscopy, and over 170 hours of MIRI imaging. This is the most intensively observed JWST field yet, designed to probe the redshift frontier beyond .
Figure 9 shows an small portion of the JOF imaging data from program 1210 to highlight the opportunities of the deep NIRCam medium-band data, here F335M and F410M. Strong emission lines are revealed easily in these images, with beautiful morphological detail. This reach will be greatly extended in program 3215 with 5 new medium bands spanning 1.6 to 3.0 .
The location of the JOF in close proximity to the HUDF and the Deep Prime JADES NIRCam imaging mosaic adds considerable value to the GOODS-S region. Wider field telescopes can capitalize on both regions, and the parallel opportunities will continue to appear with JWST. By construction, further NIRSpec observations of the HUDF can place further NIRCam exposure time in the JOF region; the position angle is at a period of low zodiacal background and is allowable with the micrometeorite constraint. Indeed, program 3215 is using this opportunity to conduct exceptionally deep spectroscopy (104 hrs of total exposure) in the HUDF, focusing on the physics of high-redshift galaxies. Further, observations with NIRISS or MIRI on the JOF are easily paired with NIRCam observations on and around the HUDF, within the JADES Deep NIRCam mosaic.
Accompanying this paper, we have released the reduced images and catalogs of the Year 1 JADES data in and around the JOF. These data are already among the very deepest observed with JWST NIRCam, with over 110 hrs of total exposure spread across 9 filters. These data are available at https://archive.stsci.edu/hlsp/jades and a feature-rich FITSmap viewer is provided at https://jades.idies.jhu.edu, with a version tailored for the public at https://jades.idies.jhu.edu/public. Observations of the Cycle 2 program 3215 are underway at the time of this writing (Oct. 2023) and will be immediately public.
The power of JWST to reveal the early history of galaxies with infrared imaging and spectroscopy is both transformative and inspiring. Many compelling observing programs have been performed or are underway, each with their own nuances and opportunities. We hope that the extraordinary depth and filter complement of the JADES Origins Field will provide a superb new view of galaxies in their youth.

### v2

@@ 6 Discussion
We have described the JWST observations of the JADES Origins Field, spanning the first three mission cycles. In total, the time invested in this single NIRCam footprint of 9 arcmin2 is already very large, roughly 360 open-shutter hours of exposure time across in 15 NIRCam imaging filters, 20 hours of NIRCam slitless spectroscopy, and over 170 hours of MIRI imaging. This is the most intensively observed JWST field yet, designed to probe the redshift frontier beyond .
Figure 9 shows a small portion of the JOF imaging data from program 1210 to highlight the opportunities of the deep NIRCam medium-band data, here F335M and F410M. Strong emission lines are revealed easily in these images, with beautiful morphological detail. This reach was greatly extended in program 3215 with 5 new medium bands spanning 1.6 to 3.0 .
The location of the JOF in close proximity to the HUDF and the Deep Prime JADES NIRCam imaging mosaic adds considerable value to the GOODS-S region. Wider field telescopes can capitalize on both regions, and the parallel opportunities will continue to appear with JWST. By construction, further NIRSpec observations of the HUDF can place further NIRCam exposure time in the JOF region; the position angle is at a period of low zodiacal background and is allowable with the micrometeorite constraint. Indeed, program 3215 used this opportunity to conduct exceptionally deep spectroscopy (91 hrs of total exposure) in the HUDF, focusing on the physics of high-redshift galaxies. Further, observations with NIRISS or MIRI on the JOF are easily paired with NIRCam observations on and around the HUDF, within the JADES Deep NIRCam mosaic.
Accompanying this paper, we have released the reduced images and catalogs of the Year 1 JADES data in and around the JOF. These data are already among the very deepest observed with JWST NIRCam, with over 110 hrs of total exposure spread across 9 filters. These data are available at https://archive.stsci.edu/hlsp/jades and a feature-rich FITSmap viewer is provided at https://jades.idies.jhu.edu, with a version tailored for the public at https://jades.idies.jhu.edu/public.
The power of JWST to reveal the early history of galaxies with infrared imaging and spectroscopy is both transformative and inspiring. Many compelling observing programs have been performed or are underway, each with their own nuances and opportunities. We hope that the extraordinary depth and filter complement of the JADES Origins Field provides a superb new view of galaxies in their youth.